OpenStudy (anonymous):

HELP! A simplified version of sin2 θ (1 + cot2 θ) = 1 is

OpenStudy (anonymous):

hmm, dunno beside just distributing

OpenStudy (anonymous):

dint see a trig sub

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

trig sub?

OpenStudy (anonymous):

i get just \[\sin^2\theta + cox^2 \theta\]

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

Whats that Tyler?

OpenStudy (anonymous):

whats the 2nd part of it

OpenStudy (anonymous):

Oh wait actually

OpenStudy (anonymous):

divide sin^2

OpenStudy (anonymous):

I did \[\sin^2\theta + \sin^2\theta*\frac{\cos^2\theta}{\sin^2\theta}\]

OpenStudy (anonymous):

sin^2 θ (csc^2 θ) ? is that what you meant?

OpenStudy (anonymous):

then you have 1+cos^2=1/sin^2

OpenStudy (anonymous):

which is -sin^2=1/sin^2

OpenStudy (anonymous):

so then you multiply sin^2 back to have -sin^4 = 1

OpenStudy (anonymous):

ALl my answers start with sin2 θ

OpenStudy (anonymous):

Then you get sin^2 + cos^2 which is 1

OpenStudy (anonymous):

sin^2 θ (1 + cot^2 θ) = 1 sin^2 θ + cos^2 θ = 1

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

??

OpenStudy (anonymous):

\[Answer = \sin^2\theta + \cos^2 \theta\]